OpenStudy (anonymous):

Solve: x divided by four equals eight 4 -4 32 -32

OpenStudy (yanasidlinskiy):

Like this???

OpenStudy (anonymous):

x/4=8 Multiply both sides by 4 to get x alone x/4(4)=8(4) x=32

OpenStudy (anonymous):

Does that make sense? :) @Antonio12345